1. A population consists of the values (2, 4, 7). List down all the possible samples of size 2 which can be drawn with replacement from this population.






1
Expert's answer
2022-05-26T06:48:51-0400

The number of possible samples which can be selected with replacement is

"N^n=3^2=9."

All the possible samples of sizes n=2 wich can be drawn with replacement from the population:

"\{(2,2),(2,4),(2,7),(4,2),(4,4),
\\(4,7),(7,2),(7,4),(7,7)\}."